The area of a rectangle is 64 square centimeters.Its perimeter is 40 centimeters.What are its dimensions?
Answer by ankor@dixie-net.com(22740)

The area of a rectangle is 64 square centimeters.
L * W = 64
:
Its perimeter is 40 centimeters.
2L + 2W = 40
simplify, divide by 2
L + W = 20
L = (20-W)
:
In the area equation, replace L with (20-W)
W(20-W) = 64
20W - W^2 = 64
arrange as a quadratic equation on the right
0 = W^2 - 20W + 64
Factors to
(W-4)(w-16) = 0
Two solutions
W = 4, then L = 16
or
W = 16, then L = 4
;
:
Check the perimeter: 2(16) + 2(4) = 40